A trapezoidal channel with side slopes 1:1 has to be designed to convey 10 m^3/sec at a velocity of 2 m/sec, so that the amount of concrete lining for the bed and sides is the minimum.
a) calculate the area of lining required for one meter length of the canal.
b) if manning's roughness coefficient n = 0.015 , calculate the bed slope of the canal
Answer #1

Lolution Data: 0-10 m/s v=am/s Area Discharge Q = 105m² velocity v 2 -- 1 A : 5m² for most economical trapezoidal section, HaArea of lining requined for one meter length of canal = (sletted perimeter) & (length of canal P= 1.369+ (2x1654), 7+1 = 6:04²i= A 0.964 i Bed slope, i- 0.9646² dlhere, es chezys con stent.